OpenStudy (anonymous):

Solve. Clear the fractions first. 3/4 + 1/5x = 1

OpenStudy (anonymous):

(3(5x)+4) / 4(5x) = 1 (15x+1)/ 20x = 1 15x+1 = 20x 1 = 5x x = 1/5